OpenStudy (anonymous):

What is the solution to the rational equation x /2x+1 -1/4= 2/2x+1

OpenStudy (saifoo.khan):

where'd you left? :O

OpenStudy (anonymous):

no i thaught u left i couldnt find u

OpenStudy (saifoo.khan):

I'm online there. LOL.

OpenStudy (anonymous):

ok on fb

OpenStudy (saifoo.khan):

Righto.

OpenStudy (anonymous):

Bring 2/2x_1 over to get \[(x-2)/2x+1=1/4\] Use cross multiplication to get \[4x-8=2x+1\]\[2x=9\]\[x=9/2\]